OpenStudy (anonymous):

secAsinA / tanA + cotA = sin^2A

OpenStudy (alekos):

I dont think theres a solution

OpenStudy (anonymous):

secAsinA / (tanA + cotA) =secAsinA /( sinA/cosA + cosA/sinA) = secAsinA( (sin^2A+cos^2A)/(sinAcosA) ) =secAsinA* (sinAcosA)/ (sin^2A+cos^2A) = sinA*sinA =sin^2A [ as secA cosA=1 and sin^2A+cos^2A=1 ]

OpenStudy (alekos):

Thanks matricked. i thought we were solving for A, but this a a prove LHS=RHS problem. no thanks to honai

OpenStudy (anonymous):

yw